4 divided x-3 minus 24 divided by x squared - 9 equals 1
Answer by jorel1380(3719)

4/(x-3)-24/(x2-9)=1
4(x+3)-24=x2-9
4x+12-24+9=x2
4x-3=x2
x2-4x+3=0
(x-1)(x-3)=0
x=1 or 3
Since it can't be 3 (4/0), the answer must be 1..